i. equilibrium condition Y= C + I
ii. Marginal Propensity To Save: 0.25
iii. Autonomous Consumption: $ 50,000
iv. Autonomous Investment: $ 25,000

a. What is the equation for the AE function?
b. Solve for equilibrium GDP.
c. What is the value of consumption when the economy is in equilibrium
d. Suppose autonomous consumption increases by $10,000 what is the new level of equilibrium?
e. What is the effect of a decrease in interest rates on equilibrium GDP? (no calculations)
1
Expert's answer
2014-10-02T13:07:48-0400
Y = C + I, s = 0.25, Ca = $50,000, I = $25,000.
a. c = 1 - s = 0.75, C = Ca + c*Y = 50,000 + 0.75Y, so the equation for the AE function will be:
AE = 50,000 + 0.75Y + 25,000 = 75,000 + 0.75Y
b. Equilibrium GDP formula is: Ye = (1/(1-c))*(Ca + I) = 4*(50,000 + 25,000) = $300,000
c. The value of consumption when the economy is in equilibrium will be:
C = Y - I = 300,000 - 25,000 = $275,000
d. If autonomous consumption increases by $10,000, the new level of equilibrium will be:
Ye = (1/(1-c))*(Ca + I) = 4*(60,000 + 25,000) = $340,000
e. A decrease in interest rates will increase equilibrium GDP, because both consumption and investment will increase.